# Intuitionism with Truth Tables: A Decision Procedure for IPL Based on RNmatrices

### v2

@@ 3 Kearns’s level valuations
(Level valuations for S4)
Assume that $Val(\mathcal{M}_{S4})$ is the set of valuations over $\mathcal{M}_{S4}$. We define the level $\mathcal{L}^{S4}_{k}$, where $k\in\mathbb{N}$, as follows:
@@
    $\mathcal{L}^{S4}_{k+1}=\{v\in\mathcal{L}^{S4}_{k}\mid\forall\alpha\in For(\Sigma_{b}),\mathcal{L}^{S4}_{k}(\alpha)\subseteq\mathcal{D}\Rightarrow v(\alpha)=2\}$
where $\mathcal{L}^{S4}_{k}(\alpha)=\{v(\alpha)\mid v\in\mathcal{L}^{S4}_{k}\}$.
@@
The set of level valuations in $\mathcal{M}_{S4}$ is defined as the intersection of every level valuation:
$\mathcal{L}_{S4}=\cap^{\infty}_{k\geq 0}\mathcal{L}^{S4}_{k}$
@@
Let $\mathcal{M}_{S4}$ be the nmatrix defined in Definition 2.4. The RNmatrix for S4 is the pair $\mathcal{R}(\mathcal{M}_{S4})=\langle\mathcal{M}_{S4},\mathcal{L}_{S4}\rangle$.
@@
Theorem 3.4 ([4]).
For every $\Gamma$ and $\varphi$: $\Gamma\vdash_{S4}^{b}\varphi$ iff $\Gamma\models_{\mathcal{R}(\mathcal{M}_{S4})}\varphi$.
3.1 Grätz’s decision procedure for S4
While the soundness and completeness of level valuations for S4 (with the rule of necessitation) were already known, it was only recently that an effective decision procedure was discovered.
Before that, J. V. Ivlev seems to have suggested the following approach [4]: remove every partial valuation $v_{p}$ whenever (1) $v_{p}(\Box A)\not\in D$ and (2) $w_{p}(A)\in D$ for every partial valuation $w_{p}$. The reason for this method is very intuitive if we consider the definition of level valuations, and indeed, it works for $2$-valued semantics, as well as for some examples in $\mathcal{M}_{S4}$, such as $\Box\alpha$ where $\alpha$ is any propositional tautology. However, this criterion is not sufficient for $\mathcal{M}_{S4}$ because it leaves behind invalid partial level valuations (consider the valid formula $\Box p\rightarrow\Box\neg\neg p$ in $\mathcal{M}_{S4}$).
In [4], Grätz identified the need to find a suitable method by refining the concept of partial valuation. To address this, he suggests a new definition for partial level valuations:
@@ 3.1 Grätz’s decision procedure for S4
[…]
  • if $\alpha\to\beta\in\Lambda$ then $\tilde{v}_{p}(\alpha\to\beta)\in{\to}^{\mathcal{A}_{S4}}(\tilde{v}_{p}(\alpha),\tilde{v}_{p}(\beta))$.
@@
A partial valuation $v_{p}:\Lambda\rightarrow\{0,1,2\}$ in $\mathcal{M}_{S4}$ is a partial level valuation in $\mathcal{R}(\mathcal{M}_{S4})$ ($v_{p}\in PLV$) iff
@@
In [4, p. 15] it is observed that, being $\Lambda$ finite, the set of $PLV$s with domain $\Lambda$ is well-defined (that is, it is not cyclic), given that the search space is finite: the set of all the partial valuations with domain $\Lambda$ is finite. Moreover, it is proved in [4] that this notion is analytic in $\mathcal{M}_{S4}$ (i.e, every partial level valuation in $\mathcal{M}_{S4}$ can be extended to a level valuation in $\mathcal{M}_{S4}$) and co-analytic (i.e, for any level valuation $v$ in $\mathcal{M}_{S4}$, the restriction of $v$ to a partial level valuation $v_{p}$ defined over a finite set $\Lambda$ closed under subformulas is a partial level valuation).
@@
$\mathcal{R}(\mathcal{M}_{S4})$ is analytic and co-analytic with PLV.
Using this, one can easily prove that the truth table method is sound and complete, that is,
Theorem 3.9 ([4]).
For every $\varphi$: $\models_{\mathcal{R}(\mathcal{M}_{S4})}\varphi$ if, and only if,for every partial valuation $v_{p}$ in $\mathcal{M}_{S4}$ defined over the set $\Lambda$ of subformulas of $\varphi$: if $v_{p}\in PLV$, then $v_{p}(\varphi)\in\mathcal{D}$.
Corollary 3.9.1 ([4]).
For every $\varphi$: $\vdash_{S4}^{b}\varphi$ if, and only if, for every partial valuation $v_{p}$ in $\mathcal{M}_{S4}$ defined over the set $\Lambda$ of subformulas of $\varphi$: if $v_{p}\in PLV$ then $v_{p}(\alpha)\in\mathcal{D}$.
@@
It is an immediate consequence of Theorems 3.9 and 3.4. ∎
Therefore, to construct a truth table, compute every partial valuation and remove any non-PLV. And this algorithmic process is a sound and complete decision procedure for S4. Obviously the soundness and completeness Theorem 3.9 (and so, the decision procedure) can be extended to inferences in S4 from a finite set of premises, taking into account that $\vdash_{S4}^{b}$ (and so $\models_{\mathcal{R}(\mathcal{M}_{S4})}$) satisfies the deduction metatheorem.
3.2 Reducing Grätz’s RNmatrix and algorithm for S4
@@
[…]
$\neg^{\mathcal{A}_{S4}}$
@@ 2 $\{0\}$$\{1\}$$\{2\}$
Of course this move requires adjusting the soundness and completeness proofs of Grätz to the new RNmatrix defined over signature $\Sigma$. Let $\vdash_{S4}$ be the consequence relation of the standard Hilbert calculus $\mathcal{H}_{S4}$ for S4 over $\Sigma$. The notion $\Gamma\vdash_{S4}\varphi$ of derivations from premises in $\mathcal{H}_{S4}$ is defined, as in the case of $\mathcal{H}^{b}_{S4}$, in terms of theoremhood.
Let $Val(\mathcal{M}^{\prime}_{S4})$ is the set of valuations over $\mathcal{M}^{\prime}_{S4}$. Consider the family of level $\mathcal{L}^{\prime S4}_{k}$ ($k\geq 0$) as in Definition 3.1, but now starting from ${\mathcal{L}^{\prime}}_{0}^{S4}=Val(\mathcal{M}^{\prime}_{S4})$ in item i), and considering formulas $\alpha\in For(\Sigma)$ in item ii). Let $\mathcal{L}^{\prime}_{S4}=\cap^{\infty}_{k\geq 0}\mathcal{L}^{\prime S4}_{k}$ and $\mathcal{R}(\mathcal{M}^{\prime}_{S4})=\langle\mathcal{M}^{\prime}_{S4},\mathcal{L}^{\prime}_{S4}\rangle$.
@@
It coincides with the proof for $\mathcal{R}(\mathcal{M}_{S4})$ found in [4, Theorems 3.4 and 3.5]. The only detail to be checked is that any axiom of $\mathcal{H_{S4}}$ is still valid in the reduced Nmatrix $\mathcal{M}^{\prime}_{S4}$. Details are left to the reader. ∎
@@
For completeness, recall that, given a Tarskian and finitary logic L and a formula $\varphi$, a set $\Delta$ of formulas is said to be $\varphi$-saturated in L if $\Delta\nvdash_{\bf L}\varphi$, but $\Delta,\alpha\vdash_{\bf L}\varphi$, whenever $\alpha\notin\Delta$. It is a well-known result that, if $\Gamma\nvdash_{\bf L}\varphi$, there exists a $\varphi$-saturated set $\Delta$ in L such that $\Gamma\subseteq\Delta$. If $\Delta$ is $\varphi$-saturated in L then it is closed, that is: $\Delta\vdash_{\bf L}\alpha$ iff $\alpha\in\Delta$. In particular, this holds in the logic S4, which is generated by $\mathcal{H_{S4}}$ over $\Sigma$. Since it contains classical logic, any $\varphi$-saturated set $\Delta$ satisfies: $\alpha\vee\beta\in\Delta$ iff $\alpha\in\Delta$ or $\beta\in\Delta$; $\alpha\land\beta\in\Delta$ iff $\alpha\in\Delta$ and $\beta\in\Delta$; and $\alpha\to\beta\in\Delta$ iff either $\alpha\notin\Delta$ or $\beta\in\Delta$. The following result is an adaptation of Lemma 3.7 in [4] to the reduced Nmatrix $\mathcal{M}^{\prime}_{S4}$:
@@
[…]
@@
Disjunction: Suppose that $v_{\Delta}(\alpha\vee\beta)=2$. Then, $\Box(\alpha\vee\beta)\in\Delta$ and so $\alpha\vee\beta\in\Delta$, by axiom T: $\Box\psi\to\psi$ and the fact that $\Delta$ is closed under logical inferences in $\mathcal{H_{S4}}$. This implies that $\alpha\in\Delta$ or $\beta\in\Delta$, hence $v_{\Delta}(\alpha)\in\mathcal{D}$ or $v_{\Delta}(\beta)\in\mathcal{D}$. This implies that $v_{\Delta}(\alpha\vee\beta)=2\in\vee^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Suppose now that $v_{\Delta}(\alpha\vee\beta)=1$. Then, $\Box(\alpha\vee\beta)\notin\Delta$ but $\alpha\vee\beta\in\Delta$ and so $\alpha\in\Delta$ or $\beta\in\Delta$. Suppose that $\Box\alpha\in\Delta$ or $\Box\beta\in\Delta$. Then, $\Box\alpha\vee\Box\beta\in\Delta$, and so $\Box(\alpha\vee\beta)\in\Delta$, given that $(\Box\alpha\vee\Box\beta)\to\Box(\alpha\vee\beta)$ is a theorem in $\mathcal{H_{S4}}$. But this is a contradiction, therefore $\Box\alpha\notin\Delta$ and $\Box\beta\notin\Delta$. This means that $v_{\Delta}(\alpha)=1$ and $v_{\Delta}(\beta)\in\{0,1\}$, or vice versa. Then, $v_{\Delta}(\alpha\vee\beta)=1\in\{1,2\}=\vee^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Finally, if $v_{\Delta}(\alpha\vee\beta)=0$ then $\alpha\vee\beta\notin\Delta$, hence $\alpha\notin\Delta$ and $\beta\notin\Delta$. This means that $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\vee\beta)=0\in\{0\}=\vee^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.

### v4

@@ 3 Gödel’s uncharacterizability theorem updated
Consider the signature $\Omega=(\{\neg\},\{\land,\lor,\rightarrow\})$ for intuitionistic propositional logic. The set of intuitionistic formulas will be denoted by $For(\Omega)$.
Definition 3.2 (Gödel [10]).
For $n\geq 1$ let $G_{n}$ be the formula expressing that “there are at most $n$ truth-values”:
$G_{n}:=\bigvee_{1\leq i<j\leq n+1}(p_{i}\to p_{j})\land(p_{j}\to p_{i}).$
Lemma 3.3 (Gödel [10]).
None of the formulas $G_{n}$ is provable in IPL.
Lemma 3.4.
If $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ is a finite Nmatrix such that $\mathcal{V}$ has exactly $n\geq 1$ elements and $\mathcal{M}$ models IPL, then the formula $G_{n}$ is valid in $\mathcal{M}$.
Proof.
For each connective $\in\Omega$ let us denote $\mathcal{O}(\#)$ by $\bar{\#}$. Since $\mathcal{M}$ models IPL then, for every $a,b\in\mathcal{V}$: (i) $a\,\bar{\vee}\,b\subseteq\mathcal{D}$ if either $a\in\mathcal{D}$ or $b\in\mathcal{D}$; (ii) $a\,\bar{\land}\,b\subseteq\mathcal{D}$ if both $a\in\mathcal{D}$ and $b\in\mathcal{D}$; and (iii) $a\,\bar{\to}\,a\subseteq\mathcal{D}$ (since $\varphi\to\varphi$ is a theorem of IPL). Given that $\mathcal{V}$ has exactly $n\geq 1$ elements, if $v$ is a valuation over $\mathcal{M}$ then, by the pigeonhole principle, there exists $1\leq i<j\leq n+1$ such that $v(p_{i})=v(p_{j})$. Let $a:=v(p_{i})$. By consideration (iii) above, $v(p_{i}\to p_{j})\in a\,\bar{\to}\,a\subseteq\mathcal{D}$, hence $v(p_{i}\to p_{j})\in\mathcal{D}$. Analogously, $v(p_{j}\to p_{i})\in\mathcal{D}$ and so, by (ii), $v((p_{i}\to p_{j})\land(p_{j}\to p_{i}))\in\mathcal{D}$. Thus, from the fact that $G_{n}$ is a disjunction where $(p_{i}\to p_{j})\land(p_{j}\to p_{i})$ is one of its disjuncts, repeated application of (i) implies $v(G_{n})\in\mathcal{D}$. ∎
Theorem 3.5.
The logic IPL cannot be characterized by a single finite-valued Nmatrix.
Proof.
Suppose that $\mathcal{M}$ is a finite Nmatrix with exactly $n\geq 1$ elements such that $\mathcal{M}$ models IPL. Suppose, by absurdum, that $\mathcal{M}$ characterizes IPL. Hence, every formula valid in $\mathcal{M}$ can be proved in IPL. By Lemma 3.4, the formula $G_{n}$ is valid in $\mathcal{M}$. By hypothesis, $G_{n}$ must be provable in IPL. But this contradicts Lemma 3.3. This shows that no finite Nmatrix $\mathcal{M}$ can characterize IPL. ∎
4 Kearns’s characterization of S4 by RNmatrices
Regarding semantics for normal modal logic, Kripke-style semantics are the usual choice. However, there are other attractive alternatives, such as algebraic semantics. In the 1980s, Kearns proposed in [19] an RNmatrix-based semantics for modal logics KT, S4 and S5. He introduced the level valuations as the set $\mathcal{F}$ of valuations to be used in his RNmatrix. According to him,
[It] is simpler than the standard [Kripke’s] account in virtue of having dispensed with possible worlds and their relations. I also think that my account is philosophically preferable to the standard account for having done this. For I do not think there are such things as possible worlds, or even that they constitute a useful fiction. (Kearns, [19, p. 86])
Figure 1 reconstructs Kearns’s Nmatrix for S4 over the signature $\Sigma_{Ke}=(\{\Box,\neg\},\{\vee\})$, where the truth-values are: $T$ (necessarily true), $t$ (contingently true), $f$ (contingently false) and $F$ (necessarily false). In this Nmatrix, the set of designated values is given by $\mathcal{D}=\{T,t\}$.
Figure 1: Tables for $\mathcal{M}_{Ke}$.
Kearns proved that, in order to fully characterize S4 using $\mathcal{M}_{Ke}$, it is sufficient to introduce the notion of level valuations, which are defined as follows:
Definition 4.1.
(Kearns’s level valuations for S4)
Let $Val^{nd}(\mathcal{M}_{Ke})$ be the set of valuations over $\mathcal{M}_{Ke}$. We define the set of $n$th-level valuations $\mathcal{L}^{Ke}_{k}$, where $n\in\omega$, as follows:
@@ 4 Kearns’s characterization of S4 by RNmatrices
    $\mathcal{L}^{Ke}_{n+1}=\{v\in\mathcal{L}^{Ke}_{n}\mid\forall\alpha\in For(\Sigma_{Ke}),\mathcal{L}^{Ke}_{n}(\alpha)\subseteq\{T,t\}\Rightarrow v(\alpha)=T\}$
where $\mathcal{L}^{Ke}_{n}(\alpha)=\{v(\alpha)\mid v\in\mathcal{L}^{Ke}_{n}\}$.
The set of level valuations in $\mathcal{M}_{Ke}$ is the intersection of the sets $\mathcal{L}^{Ke}_{n}$, for $n\geq 0$:
$\mathcal{L}_{Ke}=\bigcap^{\infty}_{n\geq 0}\mathcal{L}^{Ke}_{n}.$
The way in which level valuations work can be appreciated in Figure 2.
Figure 2: Assume that $Val^{nd}(\mathcal{M}_{Ke})=\{v_{0},v_{1},\ldots,v_{m},v_{n},v_{k},\ldots\}$. We remove a valuation $v$ in level $\mathcal{L}_{n+1}$ iff there is some $\alpha$ such that $v(\alpha)=t$ and $w(\alpha)\in\{T,t\}$ for every $w\in\mathcal{L}_{n}$.
Definition 4.2 (Kearns’s RNmatrix for S4).
The RNmatrix for S4 defined by Kearns is the pair $\mathcal{R}(\mathcal{M}_{Ke})=\langle\mathcal{M}_{Ke},\mathcal{L}_{Ke}\rangle$.
Theorem 4.3 ([19]).
For every $\Gamma$ and $\varphi$ over $\Sigma_{Ke}$: $\Gamma\vdash_{S4}\varphi$ iff $\Gamma\models_{\mathcal{R}(\mathcal{M}_{Ke})}\varphi$.
5 Grätz’s decision procedure for S4
Kearns proved that his Nmatrix for S4 with level valuations is sound and complete regarding the logic S4. However, as Section 4 explains, level valuations are defined as the intersection of an infinite family of subsets of valuation functions. How do we reproduce this procedure using finite steps? While it is true that level valuation semantics characterizes S4, it is not straightforward to decide the validity of a given formula using Kearns’s tables.
In particular, given any theorem $\varphi$, the immediate table for $\Box\varphi$ will always present a partial valuation $v$ such that $v(\Box\varphi)$ is not a designated value, thus invalidating the rule of necessitation. Grätz noticed that the problem lies in the definition of partial valuation: such partial valuations are precisely the ones that cannot be extended to full level valuations; thus, they should be ruled out by some appropriate definition.
Grätz’s Nmatrix for S4:
Instead of $4$ truth-values, Grätz use a set with three values, $0,1$ and $2$, where each one has the following intended meaning:
$v(\varphi)=0$
means that $\varphi$ is false;
$v(\varphi)=1$
means that $\varphi$ is contingently true, i.e, $\varphi$ is true but possibly false;
$v(\varphi)=2$
means that $\varphi$ is necessarily true.
From this interpretation, defining the subset $\mathcal{D}=\{2,1\}$ as the set of designated values is natural. As primitive operators, it is used negation ($\neg$), implication ($\to$), and box ($\Box$). The Nmatrix $\mathcal{M}_{S4}$ over $\Sigma_{b}=(\{\neg,\Box\},\{\rightarrow\})$ is defined as in Example 2.6 (see also Table 1).
To understand the Nmatrix for the modal operator $\Box$, we must recall the axioms that define S4. In particular, the axioms $\Box\varphi\to\varphi$ (reflexivity) and $\Box\varphi\to\Box\Box\varphi$ (transitivity). Consider the contrapositive of the reflexivity axiom: $\neg\varphi\to\neg\Box\varphi$. This means that if $\varphi$ is false, then $\Box\varphi$ is also false. The Nmatrix for $\Box$ reflects this behavior when $v(\varphi)=0$. Next, when $v(\varphi)=2$ (indicating that $\varphi$ is necessary, or in other words, $\Box\varphi$ is true), we have $v(\Box\varphi)=\{2\}$. This signifies that $\Box\varphi$ is also necessary, which implies that $\Box\Box\varphi$ is true as well. Finally, when $v(\varphi)=1$, it indicates that $\varphi$ is contingent, meaning it is possibly false, hence $\Box\varphi$ is false.
@@ 5 Grätz’s decision procedure for S4
(Grätz’s level valuations for S4)
Let $Val^{nd}(\mathcal{M}_{S4})$ be the set of valuations over $\mathcal{M}_{S4}$. We define the set of $n$th-level valuations $\mathcal{L}^{S4}_{k}$, where $n\in\omega$, as follows:
  1. i)
    $\mathcal{L}_{0}^{S4}=Val^{nd}(\mathcal{M}_{S4})$;
  2. ii)
    $\mathcal{L}^{S4}_{n+1}=\{v\in\mathcal{L}^{S4}_{n}\mid\forall\alpha\in For(\Sigma_{b}),\mathcal{L}^{S4}_{n}(\alpha)\subseteq\{1,2\}\Rightarrow v(\alpha)=2\}$
where $\mathcal{L}^{S4}_{n}(\alpha)=\{v(\alpha)\mid v\in\mathcal{L}^{S4}_{n}\}$.
The set of level valuations in $\mathcal{M}_{S4}$ is the intersection of the sets $\mathcal{L}^{S4}_{n}$, for $n\geq 0$:
$\mathcal{L}_{S4}=\bigcap^{\infty}_{n\geq 0}\mathcal{L}^{S4}_{n}.$
@@
The RNmatrix for S4 defined by Grätz is the pair $\mathcal{R}(\mathcal{M}_{S4})=\langle\mathcal{M}_{S4},\mathcal{L}_{S4}\rangle$.
@@
Theorem 5.3.
([12, Theorems 3.5 and 3.10]) For every $\Gamma$ and $\varphi$ over $\Sigma_{b}$: $\Gamma\vdash_{S4}^{b}\varphi$ iff $\Gamma\models_{\mathcal{R}(\mathcal{M}_{S4})}\varphi$.
In order to obtain a decision procedure for S4 based on the $3$-valued RNmatrix $\mathcal{R}(\mathcal{M}_{S4})$, Grätz proposes the notion of partial level valuation. The idea is as follows: given a formula $\varphi$ to be checked as a tautology in S4, construct the (finite) table generated by the Nmatrix $\mathcal{M}_{S4}$ for $\varphi$ and all of its subformulas. Then, by means of a decision procedure, it is possible to determine which rows of the table are allowed (since they correspond to level valuations); the non-allowed ones correspond to valuations over $\mathcal{M}_{S4}$ which do not correspond to level valuations, hence they are deleted. If $\varphi$ only receives a designated value in the allowed rows, then it is declared to be valid in S4, otherwise it is declared to be non-valid. Formally:
@@
  • if $#\in\{\neg,\Box\}$ and $\alpha\in\Lambda$ then $\tilde{v}_{p}(\#\alpha)\in\#^{S4}(\tilde{v}_{p}(\alpha))$;
  • if $\alpha\to\beta\in\Lambda$ then $\tilde{v}_{p}(\alpha\to\beta)\in{\to}^{S4}(\tilde{v}_{p}(\alpha),\tilde{v}_{p}(\beta))$.
@@
Let $\Lambda\subseteq For(\Sigma_{b})$ be a set of formulas closed under subformulas. A partial valuation $v_{p}:\Lambda\rightarrow\{0,1,2\}$ in $\mathcal{M}_{S4}$ is a partial level valuation over $\Lambda$ in $\mathcal{R}(\mathcal{M}_{S4})$ ($v_{p}\in PLV$) iff
@@
The set $PLV$ plays a fundamental role in the proof of soundness and completeness of the truth-tables decision procedure introduced in [12]. Some comments on this notion are in order at this point.
The existence (and uniqueness) of the set $PLV$ is the crucial fact used in his proof. In this respect, [12, p. 15] presents an algorithm (Algorithm 3.4) for computing the set $PLV$ of partial level valuations as a (constructive) truth-table. He argues that this process is well-defined since the set of all partial valuations with domain $\Lambda$ is finite, so is its powerset. Uniqueness is tacitly assumed, given the nature of the algorithm. Moreover, he proves that this notion is analytic in $\mathcal{M}_{S4}$ (i.e., every partial level valuation in $\mathcal{M}_{S4}$ can be extended to a level valuation in $\mathcal{M}_{S4}$) and co-analytic (i.e., for any level valuation $v$ in $\mathcal{M}_{S4}$, the restriction of $v$ to a partial level valuation $v_{p}$ defined over a finite set $\Lambda$ closed under subformulas is a partial level valuation). This argument, justified by the uniqueness of the set $PLV$, allows us to conclude that the truth-table procedure corresponds exactly to the level valuation semantics. A similar and fully detailed argument will be given for $\mathbf{IPL}$ in Section 9.
From the observations above, one can easily prove that the truth-table method for S4 is sound and complete, that is:
@@
([12, Theorem 3.15]) Let $\varphi$ be a formula over $\Sigma_{b}$. Then, $\models_{\mathcal{R}(\mathcal{M}_{S4})}\varphi$ if, and only if, for every partial valuation $v_{p}$ in $\mathcal{M}_{S4}$ defined over the set $\Lambda$ of subformulas of $\varphi$: if $v_{p}\in PLV$, then $v_{p}(\varphi)\in\mathcal{D}$.
Corollary 5.7.1.
For every $\varphi$ over $\Sigma_{b}$: $\vdash_{S4}^{b}\varphi$ if, and only if, for every partial valuation $v_{p}$ in $\mathcal{M}_{S4}$ defined over the set $\Lambda$ of subformulas of $\varphi$, if $v_{p}\in PLV$ then $v_{p}(\alpha)\in\mathcal{D}$.
@@
It is an immediate consequence of Theorems 5.7 and 5.3. ∎
Therefore, to construct a truth-table, compute every partial valuation and remove any non-$PLV$ in a finite number of steps. This algorithmic process is a sound and complete decision procedure for S4. Clearly, the soundness and completeness Theorem 5.7 (and hence the decision procedure) can be extended to inferences in S4 from a finite set of premises, since $\vdash_{S4}^{b}$ (and thus $\models_{\mathcal{R}(\mathcal{M}_{S4})}$) satisfies the deduction metatheorem.
6 Reducing Grätz’s RNmatrix and algorithm for S4
@@
It is easy to see that the reduced Nmatrix for S4 over signature $\Sigma$ (recall Example 2.6 and Table 1) obtained in this way is $\mathcal{M}^{\prime}_{S4}=\langle\mathcal{V},\mathcal{D},\mathcal{O}^{\prime}\rangle$, where $\mathcal{O}^{\prime}=\{\neg^{S4},\Box^{S4},\rightarrow^{\prime S4},\lor^{\prime S4},\land^{\prime S4}\}$ such that the tables for $^{\prime S4}$, for $#\in\{\to,\vee,\land\}$, are displayed in Table 2.
$\neg^{S4}$
@@ 2 $\{0\}$ $\{1\}$ $\{2\}$
Note that the Nmatrices in Table 2 were obtained by computing truth-tables, and therefore filtered sets of partial valuations. Because of this, some rows were excluded. In particular, if $v(\beta)=2$ and $v(\alpha\to\beta)=1$, we cannot find a partial valuation $w$ such that $w(\alpha\to\beta)=0$ and $w(\beta)=2$. Therefore, such a partial valuation cannot be a partial level valuation and we can safely remove it from the definition of implication. This was also discussed in [12, p. 17].
Of course considering the reduced Nmatrix over signature $\Sigma$ requires adjusting the soundness and completeness proofs of Grätz to the new RNmatrix defined over $\Sigma$. Let $\vdash_{S4}$ be the consequence relation of the standard Hilbert calculus $\mathcal{H}_{S4}$ for S4 over $\Sigma$. The notion $\Gamma\vdash_{S4}\varphi$ of derivations from premises in $\mathcal{H}_{S4}$ is defined, as in the case of $\mathcal{H}^{b}_{S4}$, in terms of theoremhood.
Let $Val(\mathcal{M}^{\prime}_{S4})$ is the set of valuations over $\mathcal{M}^{\prime}_{S4}$. Consider the family of level $\mathcal{L}^{\prime S4}_{k}$ ($k\geq 0$) as in Definition 4.1, but now starting from ${\mathcal{L}^{\prime}}_{0}^{S4}=Val(\mathcal{M}^{\prime}_{S4})$ in item i), and considering formulas $\alpha\in For(\Sigma)$ in item ii). Let $\mathcal{L}^{\prime}_{S4}=\bigcap^{\infty}_{k\geq 0}\mathcal{L}^{\prime S4}_{k}$ and $\mathcal{R}(\mathcal{M}^{\prime}_{S4})=\langle\mathcal{M}^{\prime}_{S4},\mathcal{L}^{\prime}_{S4}\rangle$.
@@
It coincides with the proof for $\mathcal{R}(\mathcal{M}_{S4})$ found in [12, Theorems 3.4 and 3.5]. The only detail to be checked is that any axiom of $\mathcal{H}_{S4}$ is still valid in the reduced Nmatrix $\mathcal{M}^{\prime}_{S4}$. Details are left to the reader. ∎
@@
The proof of completeness requires the use of a fundamental notion: $\varphi$-saturated sets.
Remark 6.3.
Recall that, given a Tarskian and finitary logic L and a formula $\varphi$, a set $\Delta$ of formulas is said to be $\varphi$-saturated in L if $\Delta\nvdash_{\bf L}\varphi$, but $\Delta,\alpha\vdash_{\bf L}\varphi$, whenever $\alpha\notin\Delta$. It is a well-known result that, if $\Gamma\nvdash_{\bf L}\varphi$, there exists a $\varphi$-saturated set $\Delta$ in L such that $\Gamma\subseteq\Delta$. If $\Delta$ is $\varphi$-saturated in L then it is closed, that is: $\Delta\vdash_{\bf L}\alpha$ iff $\alpha\in\Delta$. In particular, this holds in the logic S4, which is generated by $\mathcal{H}_{S4}$ over $\Sigma$. Since it contains classical logic, any $\varphi$-saturated set $\Delta$ satisfies: $\alpha\vee\beta\in\Delta$ iff $\alpha\in\Delta$ or $\beta\in\Delta$; $\alpha\land\beta\in\Delta$ iff $\alpha\in\Delta$ and $\beta\in\Delta$; and $\alpha\to\beta\in\Delta$ iff either $\alpha\notin\Delta$ or $\beta\in\Delta$. The following result is an adaptation of Lemma 3.7 in [12] to the reduced Nmatrix $\mathcal{M}^{\prime}_{S4}$:
@@
The unary multioperators interpreting $\neg$ and $\Box$ in $\mathcal{M}^{\prime}_{S4}$ and in the Nmatrix $\mathcal{M}_{S4}$ for S4 considered in [12] coincide, while $\to^{\prime S4}$ is different to the corresponding multioperator in $\mathcal{M}_{S4}$ (which is precisely $\to^{S4}$). In turn, the multioperators $\lor^{\prime S4}$ and $\land^{\prime S4}$ were not considered in [12, Lemma 3.7]. Thus, it is only required to prove that $v_{\Delta}(\alpha\#\beta)\in\#^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$ for every $\alpha,\beta$ and $#\in\{\lor,\land,\to\}$.
@@
Disjunction: Suppose that $v_{\Delta}(\alpha\vee\beta)=2$. Then, $\Box(\alpha\vee\beta)\in\Delta$ and so $\alpha\vee\beta\in\Delta$, by axiom T: $\Box\psi\to\psi$ and the fact that $\Delta$ is closed under logical inferences in $\mathcal{H}_{S4}$. This implies that $\alpha\in\Delta$ or $\beta\in\Delta$, hence $v_{\Delta}(\alpha)\in\mathcal{D}$ or $v_{\Delta}(\beta)\in\mathcal{D}$. This implies that $v_{\Delta}(\alpha\vee\beta)=2\in\vee^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Suppose now that $v_{\Delta}(\alpha\vee\beta)=1$. Then, $\Box(\alpha\vee\beta)\notin\Delta$ but $\alpha\vee\beta\in\Delta$ and so $\alpha\in\Delta$ or $\beta\in\Delta$. Suppose that $\Box\alpha\in\Delta$ or $\Box\beta\in\Delta$. Then, $\Box\alpha\vee\Box\beta\in\Delta$, and so $\Box(\alpha\vee\beta)\in\Delta$, given that $(\Box\alpha\vee\Box\beta)\to\Box(\alpha\vee\beta)$ is a theorem in $\mathcal{H}_{S4}$. But this is a contradiction, therefore $\Box\alpha\notin\Delta$ and $\Box\beta\notin\Delta$. This means that $v_{\Delta}(\alpha)=1$ and $v_{\Delta}(\beta)\in\{0,1\}$, or vice versa. Then, $v_{\Delta}(\alpha\vee\beta)=1\in\{1,2\}=\vee^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Finally, if $v_{\Delta}(\alpha\vee\beta)=0$ then $\alpha\vee\beta\notin\Delta$, hence $\alpha\notin\Delta$ and $\beta\notin\Delta$. This means that $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\vee\beta)=0\in\{0\}=\vee^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.